OpenStudy (anonymous):

A cellular phone company offers a contract for which the cost, C, in dollars, of t minutes of telephoning is given by C=0.25(t-500)+77.95, where it is assumed that t is greater than or equal to 500 minutes. What times will keep the cost between $126.70 and $159.95?

OpenStudy (anonymous):

\[126.70 \le.25(t-500)+77.95 \le159.95\]subtract the constant\[48.75 \le .25(t-500) \le82.00\] multiply by 4\[195 \le t-500 \le328\]add the constant\[695 \le t \le 828\]Thus, time between 695 minutes and 828 minutes will keep the costs within the stated bounds.